OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

find the slope... I suppose. So... 5/3

OpenStudy (anonymous):

and then plug in points...

OpenStudy (anonymous):

m=(-8--3)/(4-7)=(5/3) y--3=(5/3)(x-7) =y=(5/3)x-(35/3)-3 =y=(5/3)x-(44/3)